# On Sybil-Proof Mechanisms

### v4

@@ 4.1 Multiple Goods with Unit-Demand
A natural question is whether Theorem 1 extends to the case of multiple goods. With multiple goods, however, there might not even be an efficient, symmetric, Incentive Compatible, and Sybil-proof direct mechanism if some bidders demand multiple units. [^10] The VCG mechanism is for example not always Sybil-proof, see Yokoo et al., (2004). One simple setting where there are multiple non-wasteful, symmetric, Incentive Compatible, and Sybil-proof direct mechanisms is the case of unit demand bidders.
@@ 5 Conclusion
In many practical scenarios, creating Sybils is cheap but not entirely free Mazorra and Della Penna, (2023). If we assume that creating Sybils has a small constant cost for all bidders, will there be a substantially larger class of mechanisms satisfying all the desirable properties? In particular, we are interested in which distributional objectives we can achieve.
@@
Appendix A Revelation principle
Generally, subtleties arise around the players reasoning about the true number of bidders. Subsequently, we assume that bidders cannot condition their message on the number of messages the Sybil mechanism receives (because they do not know it when sending the message to the mechanism). For dominant strategy implementation case, we can then operate in a ”belief-free” model: sending message $m_{i}(v_{i})$ should be weakly-dominant for a bidder $i$ with value $v_{i}$, for all possible sets of participants in the Sybil mechanism. For the Bayesian setting bidders will also hold beliefs about the number (or more generally set of) bidders that participate in the Sybil mechanism, as we further discuss below.
A social welfare function for a variable population model is a function $f:\mathcal{D}\to\Delta(\mathbb{N})\times\mathbb{R}_{+}^{\mathbb{N}}$ with $f_{i}(v_{1},\ldots)=(0,0)$ for $v_{i}=0$ where $\mathcal{D}\subseteq c_{00}(\mathbb{R}_{+})$ is a subset of the set of $\mathbb{R}_{+}$-valued sequences with finitely many non-zero elements. The first element of $f_{i}(v)$ corresponds to the probability with which bidder $i$ gets the item, and the second element corresponds to the payment bidder $i$ makes. A social welfare function yields a direct mechanism where $\mathcal{M}=\mathbb{R}_{+}$ and $(x^{\mathcal{N}}(v),p^{\mathcal{N}}(v))=f(v|_{\mathcal{N}},0|_{\mathbb{N}\setminus\mathcal{N}})$.
[…]
  1. 1.
[…]
  2. 2.
We have the following proposition which is similar to an observation by Yokoo et al., (2004):
[…]
Proof.
[…]
  1. 1.
  2. 2.
[…]
Proof.
Appendix B Sybil-Proofness of Example 1
Proof.
We prove something slightly stronger: an agent cannot benefit from creating a Sybil and bidding arbitrarily both from the original account and the Sybil account.
We first calculate the payments from Myerson’s lemma. Let $N:=|\mathcal{N}|$. For $v\in\mathbb{R}_{+}^{N}$, we let $C:=\sum_{i=2}^{N}v_{i}$ and we have
$p_{1}(v)$ $=$ $2^{-N}\left[v_{1}\frac{v_{1}}{v_{1}+C}-\int_{0}^{v}\frac{z}{z+C}dz\right]$
$=$ $2^{-N}\left[v_{1}\frac{v_{1}}{v_{1}+C}-\ln\left(\frac{C}{v+C}\right)\right].$
Suppose that there are $n$ bidders with values $u_{1},\ldots,u_{n}$. We let $C:=\sum_{i=2}^{n}u_{i}$. If they all bid truthfully, Bidder $1$ has utility payoff
$U_{1}(u)=2^{-n}u_{1}\frac{u_{1}}{u_{1}+C}-2^{-n}\left[u_{1}\frac{u_{1}}{u_{1}+C}-\ln\left(\frac{C}{u_{1}+C}\right)\right].$
Suppose that Bidder $1$ deviates by bidding some other $y$ and creating a Sybil bidding $z$. Then his utility payoff will be
$U^{\prime}(y,z)$ $:=$ $2^{-(n+1)}u_{1}\frac{y+z}{y+z+C}$
$-2^{-(n+1)}\left[y\frac{y}{y+C}-\ln\left(\frac{C}{y+C}\right)\right]-2^{-(n+1)}\left[z\frac{z}{z+C}-\ln\left(\frac{C}{z+C}\right)\right].$
Notice that
$x\mapsto x\frac{x}{x+C}-\ln\left(\frac{C}{x+C}\right)$
is a convex function in $x$. Hence, $U^{\prime}(y,z)$ is maximized when $y=z$. We simplify $U^{\prime}(y,y)$ to
$U^{\prime}(y,y)=2^{-n}u_{1}\frac{y}{2y+C}-2^{-n}\left[y\frac{y}{y+C}-\ln\left(\frac{C}{y+C}\right)\right].$
However, this is smaller than the utility Bidder $1$ gets when bidding $y$ and not creating Sybils
$2^{-n}u_{1}\frac{y}{y+C}-2^{-n}\left[y\frac{y}{y+C}-\ln\left(\frac{C}{y+C}\right)\right],$
which is again smaller than $U_{1}(u)$ by Incentive Compatibility. Hence, Bidder $1$ has no incentive to deviate by creating Sybils. ∎
Appendix C Lemmas for the Proof of Theorem 2
@@
Let $F$ be a distribution function on $0,\infty)$, and fix $0\leq a\leq v$. Define
$U(v,a)\;=\;F(a)^{n-1}\bigl(v-a\bigr)\;+\;\int_{0}^{a}F(x)^{n-1}\,dx\;+\;(v-a)\sum_{m=1}^{n-1}\binom{n-1}{m}\,\bigl[1-F(a)\bigr]^{m}\,F(a)^{\,n-1-m}\,\frac{1}{m+1}.$
@@
First, recall the binomial identity
$\sum_{m=0}^{n-1}\binom{n-1}{m}\,\bigl[1-F(a)\bigr]^{m}\,F(a)^{\,n-1-m}\,\frac{1}{m+1}\;=\;\frac{\,1-\bigl[F(a)\bigr]^{n}\,}{\,n\,\bigl[1-F(a)\bigr]}\,.$
Hence
$\sum_{m=1}^{n-1}\binom{n-1}{m}\,\bigl[1-F(a)\bigr]^{m}\,F(a)^{\,n-1-m}\,\frac{1}{m+1}\;=\;\frac{\,1-\bigl[F(a)\bigr]^{n}\,}{\,n\,[\,1-F(a)\,]}\;-\;F(a)^{\,n-1}.$
Substitute this back into the definition of $U(v,a)$. The part in parentheses becomes
$(v-a)\,\biggl[\,\frac{\,1-F(a)^{n}\,}{\,n\,[1-F(a)]}\;-\;F(a)^{n-1}\biggr].$
Hence
$U(v,a)\;=\;F(a)^{n-1}(v-a)\;+\;\int_{0}^{a}F(x)^{\,n-1}\,dx\;+\;(v-a)\,\biggl[\frac{\,1-F(a)^{n}\,}{\,n\,[1-F(a)]}\;-\;F(a)^{n-1}\biggr].$
Factor out (va)(v-a):
$U(v,a)\;=\;\int_{0}^{a}F(x)^{\,n-1}\,dx\;+\;(v-a)\,\biggl[F(a)^{n-1}\;+\;\frac{1-F(a)^{n}}{n[\,1-F(a)\,]}\;-\;F(a)^{n-1}\biggr].$
Clearly the $F(a)^{n-1}$ terms cancel, leaving
$U(v,a)\;=\;\int_{0}^{a}F(x)^{\,n-1}\,dx\;+\;(v-a)\,\frac{\,1-F(a)^{n}\,}{\,n\,[1-F(a)]}.$
Lemma 6.
Suppose $F$ is a cdf with support $[0,1]$ twice differentiable. Let $n\geq 2$ be an integer and $k>0$. Define
$S(a)\;=\;\sum_{m=1}^{n-1}\binom{n-1}{m}\,(1-F(a))^{m}\,F(a)a^{\,n-1-m}\,\frac{k}{m+k},\quad a\in[0,1].$
Set $a_{\star}\;=\;\frac{n-2}{\,n-1\,}$. Then the function $S(a)$ is non-increasing on the interval $F^{-1}(a_{\star}),\,1]$. Moreover, if we define
$G(a)\;=\;(v-a)\,S(F(a)),\quad\text{for some }v\in[F^{-1}(a_{\star}),1],$
then $G(a)$ is also non-increasing on $F^{-1}(a_{\star}),v]$.
Since $F$ is increasing, making a change of variables, is sufficient to show it for $F(x)=x$.
First, observe that
$\frac{k}{\,m+k\,}\;=\;\int_{0}^{1}k\,x^{k-1}\,x^{m}\,dx\;=\;\int_{0}^{1}k\,x^{m+k-1}\,dx.$
Hence each term in the sum for $S(a)$ can be written as
$\binom{n-1}{m}\,(1-a)^{m}\,a^{\,n-1-m}\int_{0}^{1}k\,x^{m+k-1}\,dx.$
Interchange summation and integration to obtain
$S(a)\;=\;\int_{0}^{1}k\,x^{\,k-1}\Bigl[\sum_{m=1}^{n-1}\binom{n-1}{m}\,\bigl((1-a)\,x\bigr)^{m}\,a^{\,n-1-m}\Bigr]\,dx.$
Recognize that
$\sum_{m=0}^{n-1}\binom{n-1}{m}\,\bigl((1-a)\,x\bigr)^{m}\,a^{\,n-1-m}\;=\;\bigl[a+(1-a)\,x\bigr]^{\,n-1},$
and the $m=0$ term alone is $a^{\,n-1}$. Thus
$\sum_{m=1}^{n-1}\binom{n-1}{m}\,\bigl((1-a)\,x\bigr)^{m}\,a^{\,n-1-m}=\bigl[a+(1-a)\,x\bigr]^{n-1}-a^{n-1}.$
Therefore
$S(a)\;=\;\int_{0}^{1}k\,x^{\,k-1}\,\Bigl[\bigl(a+(1-a)\,x\bigr)^{\,n-1}-a^{n-1}\Bigr]\,dx.$
Monotonicity of $S(a)$ for $a\geq a_{\star}$. We show $S^{\prime}(a)\leq 0$ when $a\geq a_{\star}=\tfrac{\,n-2\,}{\,n-1\,}$. Differentiate under the integral:
$S^{\prime}(a)\;=\;\int_{0}^{1}k\,x^{k-1}\,\frac{d}{da}\Bigl[(a+(1-a)\,x)^{n-1}-a^{n-1}\Bigr]\,dx.$
The derivatives inside are
$\frac{d}{da}\bigl[(a+(1-a)\,x)^{n-1}\bigr]=(n-1)\,(a+(1-a)\,x)^{\,n-2}\,(1-x),\quad\frac{d}{da}\bigl[a^{n-1}\bigr]=(n-1)\,a^{\,n-2}.$
Hence
$S^{\prime}(a)\;=\;(n-1)\,k\int_{0}^{1}x^{k-1}\Bigl[(a+(1-a)\,x)^{\,n-2}\,(1-x)\;-\;a^{\,n-2}\Bigr]\,dx.$
Define
$E(a,x)\;:=\;(a+(1-a)\,x)^{\,n-2}\,(1-x)\;-\;a^{\,n-2}.$
We claim $E(a,x)\leq 0$ for all $x\in[0,1]$ whenever $a\geq\tfrac{n-2}{\,n-1\,}.$ Factor out $a^{\,n-2}$ in the first term:
$(a+(1-a)\,x)^{\,n-2}=a^{\,n-2}\Bigl[1+\tfrac{1-a}{\,a\,}x\Bigr]^{\,n-2}.$
Thus
$E(a,x)=a^{\,n-2}\Bigl[(1-x)\,\bigl(1+c\,x\bigr)^{\,n-2}-1\Bigr],\quad c=\frac{\,1-a\,}{\,a\,}.$
We can show that if $c\leq\tfrac{1}{n-2}$, then $(1-x)(1+c\,x)^{\,n-2}\leq 1$. Define
$L(x)\;=\;(1-x)\,\bigl(1+c\,x\bigr)^{n-2}.$
At the endpoints, $L(0)=1$ and $L(1)=(1-1)\,(1+c)^{n-2}=0$. A brief derivative check shows that if $c\leq\frac{1}{n-2}$, then $L^{\prime}(x)$ is non-positive on $[0,1]$, implying $L(x)$ is decreasing. Indeed, one computes
$L^{\prime}(x)\;=\;(1+c\,x)^{n-3}\Bigl[\,-\bigl(1+c\,x\bigr)\;+\;(n-2)\,c\,\bigl(1-x\bigr)\Bigr],$
and the bracketed term is never positive for $0<x\leq 1$ when $c\leq\tfrac{1}{n-2}$. Hence $L(x)$ decreases from $L(0)=1$ to $L(1)=0$, so $L(x)\leq 1$ holds on the entire interval.
$c\leq\frac{1}{n-2}$ corresponds to $a\geq\tfrac{n-2}{n-1}$. Under that condition, we get $E(a,x)\leq 0$ for all $x\in[0,1]$, and hence $S^{\prime}(a)\leq 0$. Therefore $S(a)$ is non-increasing on $\bigl[a_{\star},\,1\bigr]$.
Monotonicity of $G(a)=(v-a)\,S(a)$. Suppose $0\leq a\leq v\leq 1$ and $a\geq a_{\star}$. Define
G(a)=(va)S(a).G(a)\;=\;(v-a)\,S(a).
We have $(v-a)\geq 0$ and, from Step 2, $S(a)$ is non-increasing and nonnegative for $a\geq a_{\star}$. A short derivative check gives
$G^{\prime}(a)=\frac{d}{da}\bigl[(v-a)\,S(a)\bigr]=(v-a)\,S^{\prime}(a)\;-\;S(a).$
Since $(v-a)\geq 0$ and $S^{\prime}(a)\leq 0$, the term $(v-a)\,S^{\prime}(a)$ is non-positive. Also, $S(a)\geq 0$, so S(a)-\,S(a) is non-positive. Therefore
$G^{\prime}(a)\;\leq\;0,$
which shows $F(a)$ is non-increasing on $a_{\star},\,v]$.
For each $a\geq a_{\star}=\tfrac{n-2}{\,n-1\,}$, the function $S(a)$ is non-increasing. Furthermore, for any $v\in[a_{\star},1]$, the product (va)S(a)(v-a)S(a) is also non-increasing on $a_{\star},\,v]$.
C.1 Proof of Lemma 4
@@
Let $i$ be an agent with valuation $v$, and define $K=\{\,j\neq i\,\mid\,v_{j}\geq a\}$. Let $U(v,a)$ denote agent $i$’s expected payoff with the lottery mechanism with ticket price $a$. We want to show that if $a^{\prime}\geq a$, then $U(v,a)\geq U(v,a^{\prime})$. First, we compute $U(v,a)$. We consider two cases:
@@
$\sum_{m=1}^{n-1}\Pr(Y=m)\,\frac{v-a}{\,m+1\,}\;=\;(v-a)\sum_{m=1}^{n-1}\binom{n-1}{m}(1-F(a))^{m}F(a)^{n-1-m}\frac{1}{m+1}.$
@@
Appendix D Proof of Proposition 1
@@
Let $v_{1},\dots,v_{n-1}$ be the bids reported, and assume w.l.o.g. that they are in descending order and $n\geq m+1$. Suppose another $i=n$ agent bids $v$. The Myerson payments of agent $i$ is
$p_{i}(v,v_{-i})=v_{i}x_{i}(v,v_{-i})-\int_{0}^{v}x_{i}(t,v_{-i}),dt.$
The allocation rule is:
[…]
If $v<v_{m+1}$ then $i$’s allocation and payment is zero, so we assume that $v\geq v_{m+1}$.
@@
    In case $v>v_{1}$. If $v_{1}\geq mv_{m}$, then the payment is $v_{m}$. Otherwise, the payment is $\tfrac{m}{m+1}v_{m}+\tfrac{m}{m+1}\max\{v_{1}/m,v_{m+1}\}$.
@@
    If $v_{1}\leq mv_{m}$, the payment is $\tfrac{m}{m+1}\max\{v_{1}/m,v_{m+1}\}$.
@@
The result is trivial if $n\leq m-1$ and so we assume that $n\geq m$. A first observation is that no agent has an incentive to use Sybil bids below the largest $m+1$ bids (including his other Sybil bids) since those will just increase payments without increasing the allocation probability. A second observation is that no agent is incentivized to bid from more than two Sybils. [^13] Subsequently, we use the term ”bid from $k$ Sybils” for the case where the agent bids from $k$ accounts (his ”original” account and k1k-1 ”fake” accounts). A third Sybil will not increase the probability of getting an item, and just will increase payments. When using two Sybils, if one bids less than the m+1m+1‑st biggest bid, that Sybil will not get allocated the item, and can only increase the payment of the other Sybil, making this deviation also not profitable. Therefore, we may assume that agents bid with two Sybils with one bid larger than $v_{m}$.
We will reason by cases.
Suppose that n1n-1 agents report $v_{1},\dots,v_{n-1}$ bids. First consider the case $v<v_{1}$.
@@
    If $v_{1}\geq mv_{m}$, then the agent has no incentive to use Sybil bids, since for him the mechanism is equivalent to a uniform‑price auction.
@@
    If $v_{1}<mv_{m}$, then if he employs two Sybils, their Sybil payments are at least $\tfrac{m}{m+1}\max\{v_{1}/m,v_{m}\}$. Hence his utility with the Sybil strategy is at most
    $v-\tfrac{2}{m+1}\max\{v_{1}/m,v_{m}\}=v-\tfrac{2}{m+1}v_{m}.$
    On the other hand, if he truthfully reports his value, his utility is $\tfrac{m}{m+1}v-\tfrac{m}{m+1}\max\{v_{1}/m,v_{m+1}\}$. Observe that $v-2\tfrac{m}{m+1}v_{m}\leq\tfrac{m}{m+1}v-\tfrac{m}{m+1}v_{m}$ iff $\tfrac{1}{m+1}v\leq\tfrac{m}{m+1}(2v_{m}-v_{m+1})$, and this inequality is true since $v_{m}\geq v_{m+1}$ and $v<v_{1}\leq mv_{m}$.
Next consider the case that $v\geq v_{1}$.
@@
    If $v_{1}<m\,v_{m}$ and $v\geq m\,v_{m}$. If the agent bids truthfully, then his utility is
    $v-\tfrac{m}{m+1}v_{m}-\tfrac{m}{m+1}\max\{v_{1}/m,v_{m+1}\}\geq v-2\tfrac{m}{m+1}v_{m}.$ (6)
    Clearly, in case the agent employs two Sybils, he has no incentive to employ one with bid higher than $m\,v_{m}$ since this would allocate the item with probability $1$ but the second Sybil would increase the payment. Therefore, let’s assume that two Sybil bids are below $m\,v_{m}$. In this case, each Sybil payment is at least $\tfrac{m}{m+1}\max\{v_{1}/m,v_{m}\}$, therefore his utility in this case is at most $v-2\tfrac{m}{m+1}\max\{v_{1}/m,v_{m}\}=v-2\tfrac{m}{m+1}v_{m}$ and so, by (6), is less than or equal to his utility bidding truthfully.
Appendix E Proof of Theorem 3

### v5

@@ 6.1 Multiple Goods with Unit-Demand
A natural question is whether Theorem 1 extends to the case of multiple goods. With multiple goods, however, there might not even be an efficient, symmetric, Incentive Compatible, and Sybil-proof direct mechanism if some bidders demand multiple units. [^12] The VCG mechanism is for example not always Sybil-proof, see Yokoo et al. (2004). One simple setting where there are multiple non-wasteful, symmetric, Incentive Compatible, and Sybil-proof direct mechanisms is the case of unit demand bidders.
@@ 7 Conclusion
In many practical scenarios, creating Sybils is cheap but not entirely free Mazorra and Della Penna (2023). If we assume that creating Sybils has a small constant cost for all bidders, will there be a substantially larger class of mechanisms satisfying all the desirable properties? In particular, we are interested in which distributional objectives we can achieve.
@@
Appendix A Lemmas for the proof of Theorem 2
@@
Let $F$ be a distribution function on $0,\infty)$, and fix $0\leq a\leq v$ such that $F(a)<1$. Define
$U(v,a)\;=\;\int_{0}^{a}F(x)^{n-1}\,dx\;+\;(v-a)\sum_{m=0}^{n-1}\binom{n-1}{m}\,\bigl[1-F(a)\bigr]^{m}\,F(a)^{\,n-1-m}\,\frac{1}{m+1}.$
@@
We only need to simplify
$S\;:=\;\sum_{m=0}^{n-1}\binom{n-1}{m}(1-F(a))^{m}F(a)^{\,n-1-m}\frac{1}{m+1}.$
Use the identity $\frac{1}{m+1}\binom{n-1}{m}=\frac{1}{n}\binom{n}{m+1}$ to get
$S$ $=\frac{1}{n}\sum_{m=0}^{n-1}\binom{n}{m+1}(1-F(a))^{m}F(a)^{\,n-1-m}=\frac{1}{n}\sum_{k=1}^{n}\binom{n}{k}(1-F(a))^{k-1}F(a)^{\,n-k}$
$=\frac{1}{n(1-F(a))}\sum_{k=1}^{n}\binom{n}{k}(1-F(a))^{k}F(a)^{\,n-k}=\frac{1}{n(1-F(a))}\Bigl(1-F(a)^{n}\Bigr)=\frac{1-F(a)^{n}}{n(1-F(a))}.$
Substituting back into the definition of $U(v,a)$ yields the claim. ∎
A.1 Proof of Lemma 4
@@
Let $i$ be an agent with valuation $v$, and define $K=\{\,j\neq i\,\mid\,v_{j}\geq a\}$. Let $U(v,a)$ denote agent $i$’s expected interim payoff with the lottery mechanism with ticket price $a$. We want to show that if $a^{\prime}\geq a$, then $U(v,a)\geq U(v,a^{\prime})$. First, we compute $U(v,a)$. We consider two cases:
@@
$\sum_{m=1}^{n-1}\Pr[Y=m]\,\frac{v-a}{\,m+1\,}\;=\;(v-a)\sum_{m=1}^{n-1}\binom{n-1}{m}(1-F(a))^{m}F(a)^{n-1-m}\frac{1}{m+1}.$
@@
Appendix B Proof of Proposition 3
@@
Let $v_{1},\dots,v_{n-1}$ be the bids reported by the other agents, and assume w.l.o.g. that they are in descending order and $n\geq m+1$. Consider agent $i=n$ with value $v$ who (when using one account) submits a bid $v$. The Myerson payment of agent $i$ is
$p_{i}(v,v_{-i})=v\cdot x_{i}(v,v_{-i})-\int_{0}^{v}x_{i}(t,v_{-i})\,dt.$
Allocation rule. Example 6.1 runs the uniform-price auction iff $v_{(m+1)}<v_{(1)}/m$ (order statistics of the full profile), and otherwise runs the $(m+1)$-lottery among the top $m+1$ bids. Fixing $v_{-i}$, this yields the following interim allocation rule for bidder $i$. Let $r:=\max\{v_{m+1},v_{1}/m\}$. Then
$x_{i}(t,v_{-i})=\begin{cases}1,&\text{if }\max\{t,v_{1}\}>m\,v_{m}\text{ and }t\geq v_{m},\\[3.0pt] \frac{m}{m+1},&\text{if }\max\{t,v_{1}\}\leq m\,v_{m}\text{ and }t\geq r,\\[3.0pt] 0,&\text{otherwise.}\end{cases}$
(In particular, when $v_{m+1}\leq t<v_{m}$, the lottery branch requires $t\geq v_{1}/m$.)
If $v<r$ then $x_{i}(v,v_{-i})=0$ and $p_{i}(v,v_{-i})=0$, so we assume $v\geq r$.
Myerson payments.
@@
    If $v>v_{1}$. If $v_{1}\geq mv_{m}$, then the payment is $v_{m}$. Otherwise, the payment is $\tfrac{m}{m+1}v_{m}+\tfrac{m}{m+1}\max\{v_{1}/m,v_{m+1}\}$.
@@
    If $v_{1}\leq mv_{m}$, then (since $v\geq r$) the payment is $\tfrac{m}{m+1}\max\{v_{1}/m,v_{m+1}\}$.
@@
Sybil deviations. The result is trivial if $n\leq m$ (then everyone receives an item and pays 0), so assume $n>m$.
Consider a unit-demand agent with value $v$ who deviates by submitting $k\geq 1$ bids $b^{1},\dots,b^{k}\geq 0$ (accounts), with no restriction that any $b^{\ell}$ equals $v$. Let $b^{(1)}\geq b^{(2)}\geq\cdots\geq b^{(k)}$ be the bids sorted.
A first observation is that no agent has an incentive to use Sybil bids below the largest $m+1$ bids (including his other Sybil bids) since those will just increase payments without increasing the allocation probability. A second observation is that no agent is incentivized to bid from more than two Sybils. A third Sybil will not increase the probability of getting an item, and just will increase payments. When using two Sybils, if one bids less than the m+1m+1‑st biggest bid, that Sybil will not get allocated the item, and can only increase the payment of the other Sybil, making this deviation also not profitable. Therefore, we may assume that agents bid with two Sybils with one bid larger than $v_{m}$. Now fix a two-bid deviation $(a,b)$ with $a\geq b$.
If the deviation induces the uniform-price auction, it is dominated by a single bid. Indeed, removing the lower bid $b$ can only weakly decrease the $(m+1)$-st order statistic and leaves the top bid unchanged, so if the profile with $(a,b)$ triggers the auction branch then the profile with only $a$ also triggers the auction branch. The lower bid $b$ cannot increase the probability of receiving a unit above $1$ (unit demand) and can only (weakly) increase payments. Therefore such a deviation is weakly dominated by a single-bid deviation with bid $a$, which cannot be profitable by incentive compatibility.
Thus it remains to consider two-bid deviations that induce the lottery. In that case both bids have allocation probability $q:=\frac{m}{m+1}$. Moreover, for each of the two accounts, its opponent set contains the $m$ bids $v_{1},\dots,v_{m}\geq v_{m}$ and the other account’s bid (which is $\geq v_{m}$ in this lottery-relevant case), so the $(m+1)$-st highest opponent bid is at least $v_{m}$. From the payment formula in Case 2.1 (applied to the relevant opponent profile of each account), each account’s expected payment is at least $qv_{m}$. Hence the total expected payment across the two accounts is at least $2qv_{m}=\frac{2m}{m+1}v_{m}$, and the deviating utility is at most
$v-\frac{2m}{m+1}v_{m}.$ (6)
Comparison with truthful utility. We now show truthful bidding yields utility at least (6) in all cases; this implies no Sybil deviation is profitable.
First consider $v<v_{1}$.
@@
    If $v_{1}\geq mv_{m}$, the mechanism coincides with the uniform-price auction for the relevant range and Sybil deviations are ruled out by the dominance argument above (auction-branch deviations reduce to a single bid, which is unprofitable by IC).
@@
    If $v_{1}<mv_{m}$, then truthful utility $U$ (for a single bid) equals
    $U=\frac{m}{m+1}\Bigl(v-\max\{v_{1}/m,v_{m+1}\}\Bigr)\geq\frac{m}{m+1}(v-v_{m}),$
    since $\max\{v_{1}/m,v_{m+1}\}\leq v_{m}$ in this case. Because $v<v_{1}\leq mv_{m}$, we have $v\leq mv_{m}$, which is equivalent to
    $v-\frac{2m}{m+1}v_{m}\leq\frac{m}{m+1}(v-v_{m})\leq U.$
    Together with (6), this rules out profitable Sybil deviations.
Next consider $v\geq v_{1}$.
@@
    If $v_{1}<mv_{m}$ and $v\geq mv_{m}$, then by the payment computation in Case 1.2., truthful utility is
    $v-\tfrac{m}{m+1}v_{m}-\tfrac{m}{m+1}\max\{v_{1}/m,v_{m+1}\}\geq v-2\tfrac{m}{m+1}v_{m}.$
    Any Sybil deviation that induces the auction branch is dominated by a single-bid deviation (hence not profitable by IC), and any Sybil deviation that induces the lottery has utility bounded by (6). Thus no Sybil deviation is profitable.
Appendix C Proof of Theorem 4